OpenStudy (anonymous):

Alma was given some money as a gift and deposited it into a savings account. She makes a withdrawal of the same amount at the end of each month. At the end of the 5th month, her balance was $442. At the end of the 13th month, her balance was $226. How much money was Alma given as a gift?

OpenStudy (anonymous):

At the end of the 5th month, you have 442. At the end of the 13th, you have just 226. That means that over 8 months, you've taken out 216 bucks. (442 - 226 = 216.) So divide that by 8 to see how much comes out each month: 216 divided by 8 = 27 bucks each month. That means that for each of the months leading up to (and including the 5th), you took out 27 bucks. So that's 27 x 5, which is . . . 135. So add 135 dollars back onto the 442 you had at the end of the 5th month, and it seems you started with... 135 + 442 = 577 dollars. Double check: 577 - (27 x 5) = 442. Then 442 - (27 x 8) = 226. And 577 - (27 x 13) = 577 - 351 = 226 too.